The population standard deviation for the height of college basketball players is 3.2 inches. If we want to estimate 92% confidence interval for the population mean height of these players with a 0.8 margin of error, how many randomly selected players must be surveyed? _____ (Round up your answer to nearest whole number)
Answer #1

Answer:

The population standard deviation for the height of college basketball players is 3.2 inches.

Population standard deviation==3.2 inches

Margin of error = E = 0.8

At 92% confidence level


=1-92%=1-0.92=0.08

Sample size(n):

Therefore Sample Size n=49.